OpenStudy (joehoffy):

please help for a medal...What is the product of 3.4 × 10^−14 and 1.8 × 10^28? A. 6.12 × 10^14 B. 9.4 × 10^14 C. 5.2 × 10^42 D. 6.12 × 1042

OpenStudy (alexandervonhumboldt2):

first of all what is 3.4*1.8=?

OpenStudy (joehoffy):

6.12

OpenStudy (alexandervonhumboldt2):

yeah now use the rule: \[\Huge\bf a^b*a^c=a^{b+c}\]

OpenStudy (alexandervonhumboldt2):

thus we have 10^-14*10^28=10^{-14+28}=?

OpenStudy (joehoffy):

in which case a and b would be what?

OpenStudy (alexandervonhumboldt2):

it is the rule

OpenStudy (alexandervonhumboldt2):

it applies for all numbers

OpenStudy (joehoffy):

it gives me 10^14

OpenStudy (alexandervonhumboldt2):

yeah so we have 6.12*10^14. so which choice?

OpenStudy (joehoffy):

A!

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (joehoffy):

thank you so much you are a hero

OpenStudy (alexandervonhumboldt2):

=]